Jefferson Lab Experiment E01-001

"New Measurement of (G_E/G_M) for the Proton"
Rogues Gallery:

Spokespersons -Ralph Segel and John Arrington

Physics Liaison - Ole Hansen

Run Coordinators - Dave Gaskell and Dave Meekins

Graduate Student - Issam Qattan

The SRCoW (Super-Rosenbluth Collaborator of the Week) this week is:
Everyone who signed up for extra shifts
 "We filled four days of shifts in no time, with very little notice" (one or two spots still available).

Previous Winners:
Nature's Fury
"Because twice is enough"
Kathy McCormick
"For last minute saves: hole drilling, target training, and target cooling"
Dave Meekins
"For last minute rescues on just about everything else"
Mark Jones (JLab)
"For his exceptional effort in coordinating the software and setup for both e01001 and e01020"

NEWS (old):  The Hall will close a day early, and the cryotarget cooldown will now begin on tuesday.   As a result, we have a few new shifts on the schedule at the beginning of the run.  Please be a good collaborator and sign up for an extra shift if you will be at JLab.

NEWS: Due to a damaged transfer line, the cryotarget cooldown has been delayed.  The problem is being fixed, and the target cooldown should begin tomorrow, so it looks as though we are basically back on the original schedule.

May 5: After brief bouts with target problems, low VDC efficiency, and poor beam quality, this evening we have had high quality 50 uA beam. At 11:20 pm we finished our first data point (Q^2=2.65, E=2.262).  Hazah!!

May 7th: A CHL/ESR crash at about 2:30 pm informs us that a thunderstorm is on its way.  It is now 9pm, and we are cooling down the HRS magnets and starting the target cooldown procedure.  We expect to have the hall ready for beam sometime after midnight.

May 12th:  Excellent running conditions since beam came back on the 7th (except for a 6-hour accelerator down that we are recovering from now).  We have finished the 2-pass and 5-pass running at 1.1 GeV/pass, and should finish the 3-pass running by tuesday.  At that point, we will have 8 data point finished, with 7 more to go after the linac energy change on tuesday.  If beam availability is high, we should be able to make up for lost time and finish by the 20th (keep your fingers crossed).

May 13th:  Well, it turns out that lightning can and does strike the same experiment twice (I guess we didn't keep our fingers crossed tightly enough).  CHL and ESR are down, the target is warming, and my uninformed guess is that we will not have any beam back before the scheduled shutdown tomorrow.  If so, we will come back Tuesday night or (more likely) Wednesday at the lower beam energy.

May 14th:  More bad news, I'm afraid.  On May 7th, when work was being done on the target as part of the recovery from the CHL crash, safety windows were put over the spectrometer entrance windows (1/8" aluminum), and over the scattering chamber exit windows (1/8" Lexan).  They were not removed after the work was done, and so the 3-pass and 5-pass data (6 points) were taken with the windows in place, while the 2-pass data (2 points) were taken without the windows.  After beam comes back at the new linac energy, we will take some runs with and without the windows, and decide on the most efficient way to take data to check the effect of the windows and/or take backup data in case they cannot be adequately corrected.

May 16th: Still trying to recover from the May 13th CHL crash.  In the meantime, Kees has extended us until noon on friday (and will have MCC look into the possibility of an extension into the weekend if necessary).  With the extra time, we should be able to take some data to test the effect of the windows, and get some new high-epsilon data as a backup to the earlier 5-pass data.  This of course assumes that we get beam back fairly soon.

May 20th:  We have taken all three Q^2 points at 3-pass, and are currently switching over to 5 pass.  Things have been running smoothly for the last  couple of dyas, and we should be able to get everything we need by friday.  Thanks to everyone who signed up for the extra shifts - I was thrilled to see the schedule fill up so quickly. (A few shifts are still available).

May 23rd:  No more beam energy changes!  We are at four pass, and will stay here until the end of the run.  It took most of a shift to get the beam spot size and raster size correct, which has put us behind schedule.  However, it we have smooth running from now until the end, we should have everything done by noon on friday.

ABU and rough overall efficiency (.ps)   (just ABU/calendar time, probably ~10% low due to configuration changes)

 Contact Information

 Run Information